Reports 1-1 of 1 Clear search Modify search
MIF (General)
takaaki.yokozawa - 8:09 Saturday 24 January 2026 (36214) Print this report
Status of the increasing laser power
I checked the status of the increasing laser power.
We can succeeded to increasing the laser power up to 13 W, but the IR trans power decreased during the increasing, then locked loss happened.
Fig.1. and Fig.2. showed the trend of the suspensions, some drift detected in IMMT2 and PRM?

Before moving the next step, we need to check the OLTF for the ASC loops (sure)
Images attached to this report
Search Help
×

Warning

×